OpenStudy (anonymous):

Any help would be great.... The area of a square garden is 36x^2-60x+25. What is the perimeter of the garden if x=3 feet?

OpenStudy (anonymous):

\[ 36x^2-60x+25=(6x-5)^2\]

OpenStudy (anonymous):

so the sides are each \(6x-5\) and the perimeter is \[4(6x-5)\]

OpenStudy (anonymous):

they really reach for these word problems don't they?

OpenStudy (anonymous):

is 6x the square root of the first number or is it related to -60x?

OpenStudy (anonymous):

oh \(x=3\) so each sides is \[6\times 3-5=18-5=13\] and the perimeter is \(4\times 13\)

OpenStudy (anonymous):

okay thank you! these really are not in my strongest area for math.. thanks again!